OpenStudy (anonymous):

1000=(10000)/(5+1245e^(-0.97t) need help to solve this pretty please?

OpenStudy (badhi):

$$\begin{align*} 1000&=\frac{10000}{5+1245e^{-0.97t}}\\ 5+1245e^{-0.97t}&=10\\ 1245e^{-0.97t}&=5\\ e^{-0.97t}=\frac{1}{249}\\ -0.97t&=\ln\left[\frac{1}{249}\right]\\ t=\frac{\ln(249)}{0.97} \end{align*}$$